# PET-MAD, a universal interatomic potential for advanced materials modeling

### v1

@@ II.8 Quantum nuclear effects in NMR crystallography
Figure 7: Computed chemical shielding distribution of 1H in $\alpha$ succinic acid crystals, using PET-MAD (blue), PET-MAD-LoRA (light blue) and PET-Bespoke (orange) models for PIMD (solid) and MD (dotted) simulations. Shieldings of the static geometry (PET-MAD only) are shown as vertical lines (carboxylic H3 site around 15 ppm and an average of the aliphatic H1 and H2 sites around 28 ppm).
In fact, NMR shieldings result from the averaging of instantaneous values of structures distorted by thermal and quantum fluctuations, which can be sampled computationally by performing MLIP-driven (PI)MD simulations and then evaluating the chemical shieldings using a bespoke machine learning model for succinic acid crystals. We simulate in particular the $\alpha$ polymorph of succinic acid, one of the structures investigated in a previous study[63] from which we also obtain the training structures to build bespoke shielding and MLIP models (the test set accuracies (MAE) on energies| forces are 12.5 meV/atom|106.1 meV/Å for PET-MAD, 3.1 meV/atom|86.0 meV/Å for PET-Bespoke and 2.0 meV/atom|64.5 meV/Å for PET-MAD-LoRA). In line with the excellent accuracy of PET-MAD for molecular crystals, we observe near-perfect agreement between the distributions of chemical shieldings obtained with the three potentials (Figure 7), with both classical and quantum sampling – the latter showing a large downward shift and broadening of the 1H shielding distributions, especially for the H-bonded protons, qualitatively the same observation that Engel and coworkers make in Ref. 63 who employ a bespoke MLIP that was trained on the significantly more expensive PBE0-MBD level of theory. As in the case of water, these results demonstrate the ability of PET-MAD to describe quantum nuclear fluctuations. By combining MLIP-driven sampling with models of functional properties, such as the chemical shieldings, one can extend the reach of a general-purpose model beyond structural and energetic predictions.
@@ II.9 Dielectric response of barium titanate
Figure 8: Temperature-dependent dielectric response of BTO for three PET models. Vertical dotted lines correspond to phase transition temperatures estimated from the relative chemical potentials. The inset shows the collective variable landscape distinguishing three separate phases. The models consistently identify the different phases, as evidenced by the contour plots at 0.5 probability centered on the basins corresponding to each phase.
@@
Barium titanate (BaTiO3, BTO) is a prototypical ferroelectric perovskite that undergoes a sequence of temperature-dependent structural phase transitions: rhombohedral (R, below 183 K), orthorhombic (O, 183–278 K), tetragonal (T, 278–393 K), and cubic (C, above 393 K) [64]. The ferroelectricity in BTO arises from off-center displacements of the Ti atom within the oxygen octahedron, which break centrosymmetry in the lower-temperature phases. These displacements are fundamental to its ferroelectric behavior and phase transitions,[65] making BTO an ideal model system for studying ferroelectricity in perovskites.
@@
To evaluate the capability of PET-MAD in characterizing ferroelectricity in this material, we perform flexible-cell MD simulations for a 320-atom model of BTO over a temperature range of 40–400 K at constant ambient pressure, following the same protocol as in Ref. 66. Once again, we recompute consistent energetics for the material-specific dataset, to build bespoke and fine-tuned PET models. Test-set MAEs on energy|forces are 12.67 meV/atom |27.96 meV/Å for PET-MAD, 0.23 meV/atom |9.41 meV/Å for the bespoke model, and 0.12 meV/atom |3.92 meV/Å for the LoRA model, respectively.
@@
We then analyze the sampled configuration within a reduced-dimensionality collective variable landscape derived from a subset of neighbor density coefficients [66] (shown in the inset of Figure 8). Measuring the populations of different phases by a clustering algorithm allows us to estimate their relative chemical potential, and hence the transition temperatures between the different phases. Similar to what is observed in Ref. 66, the transition temperatures are heavily underestimated with respect to experiments (due to a combination of DFT shortcomings and finite-size effects), but there is excellent agreement between PET-MAD and the two bespoke models. The largest discrepancy, for the T-C transition, is below 30 K—a remarkable accuracy for a model with such broad applicability.
@@
To further assess the reliability of PET-MAD in predicting functional properties, we also compute the temperature-dependent static dielectric tensor $\boldsymbol{\epsilon}_{r}$ (Figure 8), by estimating the covariance of the total polarization, computed in turn using an equivariant linear model trained on the polarization dataset of Ref. 66. All models predict qualitatively the large value of $\boldsymbol{\epsilon}_{r}$ for the paraelectric, cubic phase, which increases greatly as it approaches the ferrolectric transition temperature, and then takes smaller values in the tetragonal and orthorhombic phases (which have multiple inequivalent optical axes). As it is the case of the transition temperatures, the two specialized models are in near-perfect agreement, while the PET-MAD simulations slightly underestimate $\boldsymbol{\epsilon}_{r}$ in the T phase, and yield a curve in the C phase that is shifted by about 25 K, consistent with the lower Curie temperature. Overall, this final example demonstrates once more that, despite considerably larger test errors than for a bespoke model, PET-MAD captures with close-to-quantitative accuracy subtle physical effects, and can be used to drive complicated, advanced materials modeling protocols.
@@ III Discussion
PET-MAD pushes the boundaries of what generally-applicable “universal” machine learning potentials can achieve, providing out-of-the-box semi-quantitative accuracy for both inorganic and organic materials benchmarks, as well as for six demonstrative applications covering a very broad range of classes of materials and advanced simulation techniques. Moreover, it does so while being trained on a data set that is orders of magnitudes smaller than those used for current state-of-the-art models. This observation suggests that the computational budget for training set construction can be best allocated by pursuing internal consistency and a high degree of structural and chemical diversity, two guiding principles which underlie the construction of the MAD dataset we introduce here. A compact dataset facilitates model training and optimization, as well as the transition to more accurate, and computationally demanding, approximations to the electronic-structure problem.
The PET architecture we use is not constrained to follow rigorous rotational invariance; nonetheless, it learns to make predictions that are invariant to a high degree of accuracy. It achieves a high expressiveness thanks to its transformer module, while being faster than all the conservative models we considered, despite the fact we have not yet dedicated substantial effort to optimizing its implementation. The availability of a reliable and inexpensive uncertainty quantification framework makes it possible to assess model error, and also its propagation to the ultimate quantity of interest. In cases where this error is larger than desired, a simple fine-tuning procedure can be applied to further improve the accuracy for a specific target system, while maintaining the ability to make qualitative predictions for the high-diversity test set. PET-MAD is easily accessible from several atomic-scale simulators thanks to its integration within the metatensor[67] ecosystem. It provides an efficient and accurate solution to perform simple and advanced atomistic simulations of materials, and we expect it will serve as inspiration for future developments of both datasets and model architectures, as well as a reliable engine to drive the design and discovery of new and better materials.
IV Materials and Methods
We provide a brief but complete overview of the procedure we followed to construct the MAD dataset, to train the PET-MAD model, and to fine-tune single-purpose models. The simulation protocols for the advanced simulation examples (Sections  II.4-II.9) follow closely those of the cited publications, and are briefly discussed in the ST for the convenience of the reader.
@@ IV.2 First-principles calculations
All datasets in this work (i.e. MAD and the reference subsets used for benchmarking, as well as the single-purpose dataset for the six case studies) were built by computing energies and forces using an identical density functional theory (DFT) setup. We used the Quantum Espresso v7.2 package [68], compiled with the SIRIUS[69] libraries as the primary DFT engine and AIIDA [70] as a workflow and task manager. To reduce the possibility that some simulations would converge to ambiguous magnetization states, all simulations were performed in a non-magnetic setting, using the PBEsol [71] exchange-correlation functional. Even though this choice is known to have shortcomings for several classes of materials – as shown by the discrepancy of some of our examples with experiments – it allows us to generate predictions for a highly diverse system with a high degree of stability and converged numerical parameters. We described valence and semi-core electrons using the standard solid-state pseudopotentials library (SSSP) v1.2 (efficiency set) [72], with the most stringent settings for plane-wave and charge-density cutoffs across the 85 elements (110 Ry and 1320 Ry, respectively). As discussed in the SI, using different settings to match the maximum cutoff of each structure would give substantial inconsistencies between the description of some atomic types. Electronic smearing and partial occupancies were described with a Marzari-Vanderbilt-DeVita-Payne cold smearing [73] with a spread of 0.01 Ry. In all periodic dimensions, the first Brillouin zone was sampled with a $\Gamma$-centered grid with a resolution of 0.125 $\pi$-1, while only a $\Gamma$ point was used along the non-periodic dimensions. In the case of non-periodic systems we also applied a truncation scheme of the Coulomb potential to avoid the interaction of periodic replicas of the system through the periodic boundary conditions: the Sohier-Calandra-Mauri method from Ref. 74 for 2D systems and the Martyna-Tuckerman correction [75] for 0D systems. Along each non-periodic direction we additionally applied a 25 Å vacuum to ensure convergence of the truncation methods. A compositional baseline based on isolated atom energies was subtracted from the DFT energies of each structure to improve the numerical stability of training.
For most of the MAD subsets listed in Section II.1, the DFT settings provided above ensured a good convergence rate of $>95\%$, i.e. convergence was not achieved in less than 5% of the cases. The exception is MC3D-random, for which only about 55% of the attempted simulations converged. This result does not come as a surprise and can be explained by the extremely non-equilibrium state of these structures.
@@ IV.3 PET model
As our machine learning approach for fitting interatomic potentials, we employ the Point Edge Transformer (PET) [19], which has been shown to achieve state-of-the-art performance across a diverse set of benchmarks spanning both molecular and materials domains. In essence, PET is a graph neural network (GNN) where each message-passing layer is implemented as an arbitrarily deep transformer [76]. Specifically, PET maintains feature vectors (or messages) $f_{ij}^{l}$ for every directed bond between atoms $i$ and $j$ that lie within a specified cutoff radius. Here, the superscript $l$ denotes the message-passing layer index, and each $f_{ij}^{l}$ is a fixed-size vector (a token) of dimension $d_{\mathrm{PET}}$. These intermediate representations are updated at each message-passing layer by a transformer with an adjustable number of internal layers, invoked for each atomic environment. At each message-passing layer, for each atom $i$, the input tokens to the transformer consist of the message vectors $\{f_{ji}^{l}\}_{j}$ coming from all neighbors $j$ to the central atom $i$. The transformer then applies a sequence-to-sequence transformation in a permutation-covariant manner. Its outputs are subsequently interpreted as the new set of outbound messages from atom $i$ to each neighbor $j$, $\{f_{ij}^{l+1}\}_{j}$. Geometric information (i.e., the 3D positions of neighbors) and chemical species are also incorporated into this process. To obtain the desired target property (e.g., energy), PET applies feed-forward neural networks to each representation $f_{ij}^{l}$ and sums all the outputs across bonds $ij$ and layers $l$. The PET architecture imposes no explicit rotational symmetry constraints, but learns to be equivariant through data augmentation. This unconstrained approach yields high theoretical expressivity: even a single layer of the model acts as a universal approximator featuring virtually unlimited body order and angular resolution. A detailed specification of the architecture and the full functional form can be found in Ref. 19.
@@ IV.4 Training of PET-MAD
We split the MAD dataset into into training, validation, and test parts by separately shuffling each of the eight subsets, and selecting respectively 80%, 10%, and 10% of the structures. The composition contribution to the total energy of the structures was fitted using a simple linear model and subtracted to improve the training behavior. We used a Pareto-optimal architecture (see the details on hyperparameters optimization in the Supplementary Text) with a cutoff radius of 4.5 Å, 2 message-passing layers, each containing 2 transformer layers with a token size of 256, 8 heads in the multi-head attention layer, 512 neurons in the output multi-layer perceptron. Training was performed using the PyTorch framework [77] and the metatrain package [67] on 4 NVIDIA H100 GPUs with a batch size of 22 structures per GPU for a total of 405 epochs and took about 20 hours. We used the Adam [78] optimizer with an initial learning rate (LR) of $10^{-4}$. For the first 100 epochs, we enabled LR warm-up, which increased the LR linearly from 0.0 to $10^{-4}$, and then applied an LR scheduler, which halved the LR every 250 epochs. The loss function was based on a root-mean-square error difference between the model’s energy and force predictions and the corresponding target values, with a scaling factor of 0.1 applied to the energy contribution. Mean absolute errors of the trained PET-MAD model were 7.3 meV/atom|43.2 meV/Å  in energies and forces prediction on a train set. Validation set errors were 14.7 meV/atom|72.2 meV/Å  , respectively.
@@ IV.5 Fine-tuning
Fine-tuning of the pre-trained PET-MAD model was performed using a parameter-efficient fine-tuning technique based on low-rank adaption (LoRA)  [27]. This method is widely used in the ML community, since it allows to mitigate the catastrophic forgetting—the phenomenon of losing the accuracy on the base dataset while tuning the model on the new dataset, which is inherent to the conventional fine-tuning, where all the weights of the model are trainable [79]. In the LoRA approach, the weights of the base model are frozen during the training, while an additional set of trainable weights composed of two low-rank matrices is added to each attention block of the model with a regularization factor, which controls the influence of the low-rank matrices on the model’s weights. By adjusting rank and regularization factor, one can optimize the fine-tuning workflow to achieve better performance on specific tasks, without compromising entirely the ability to perform the more general task. In this work, we used a rank of 8 and a scaling parameter of 0.5 for all LoRA fine-tuned models, unless otherwise specified.
In terms of raw accuracy on the specialized datasets, the results in Section II and Section S7 indicate that fine-tuning is always beneficial compared to training a specialized ad-hoc model in the low-data regime. On larger datasets, while in some cases a specialized model trained from scratch can exceed the accuracy of a fine-tuned model, this is not the case for many of the systems investigated in this work, including barium titanate, succinic acid, lithium thiophosphate, and high-entropy alloys.
LoRA-fine-tuned models retain varying degree of accuracy (see the Section S6 for details) on the generic structures from the MAD dataset, while still providing practically equivalent observables from an ad-hoc fully specialized model. In general, we therefore recommend to employ LoRA if fine-tuning PET-MAD for a specific application.

### v2

@@ II.8 Quantum nuclear effects in NMR crystallography
Figure 8: Computed chemical shielding distribution of 1H in $\alpha$ succinic acid crystals, using PET-MAD (black), PET-MAD-LoRA (orange) and PET-Bespoke (blue) models for PIMD (solid) and MD (dotted) simulations. Shieldings of the static geometry (PET-MAD only) are shown as vertical lines (carboxylic H3 site around 15 ppm and an average of the aliphatic H1 and H2 sites around 28 ppm).
NMR shieldings result from the averaging of instantaneous values of structures distorted by thermal and quantum fluctuations, which can be sampled computationally by performing MLIP-driven (PI)MD simulations and then evaluating the chemical shieldings using a bespoke machine learning model for succinic acid crystals. We simulate in particular the $\alpha$ polymorph of the succinic acid, one of the structures investigated in a previous study [60] from which we also obtain the training structures to build bespoke shielding and MLIP models (the test set accuracies (MAE) on energies|forces are 12.5 meV/atom|106.1 meV/Å for PET-MAD, 3.1 meV/atom|86.0 meV/Å for PET-Bespoke and 2.0 meV/atom|64.5 meV/Å for PET-MAD-LoRA). In line with the excellent accuracy of PET-MAD for molecular crystals, we observe near-perfect agreement between the distributions of chemical shieldings obtained with the three potentials (Figure 8), with both classical and quantum sampling – the latter showing a large downward shift and broadening of the 1H shielding distributions, especially for the H-bonded protons, qualitatively the same observation that Engel and coworkers make in Ref. 60 who employ a bespoke MLIP trained on PBE0-MBD reference calculations. As in the case of water, these results demonstrate the ability of PET-MAD to describe quantum nuclear fluctuations. By combining MLIP-driven sampling with models of functional properties, such as the chemical shieldings, one can extend the reach of a general-purpose model beyond structural and energetic predictions.
@@ II.9 Dielectric response of barium titanate
Figure 9: Temperature-dependent dielectric response of BTO for three PET models: the pre-trained PET-MAD (black), the bespoke PET model (blue) and a LoRA-finetuned PET-MAD (orange). Vertical dotted lines correspond to phase transition temperatures estimated from the relative chemical potentials. The inset shows the collective variable landscape distinguishing three separate phases. The models consistently identify the different phases, as evidenced by the contour plots at 0.5 probability centered on the basins corresponding to each phase.
@@
To conclude our benchmarking series, we study the ability of PET-MAD to describe the dielectric properties of materials using barium titanate system as an example. Barium titanate (BaTiO3, BTO) is a prototypical ferroelectric perovskite that undergoes a sequence of temperature-dependent structural phase transitions: rhombohedral (R, below 183 K), orthorhombic (O, 183–278 K), tetragonal (T, 278–393 K), and cubic (C, above 393 K) [61]. The ferroelectricity in BTO arises from off-center displacements of the Ti atom within the oxygen octahedron, which break centrosymmetry in the lower-temperature phases. These displacements are fundamental to its ferroelectric behavior and phase transitions,[62] making BTO an ideal model system for studying ferroelectricity in perovskites.
@@
To evaluate the capability of PET-MAD in characterizing ferroelectricity in this material, we perform flexible-cell MD simulations for a 320-atom model of BTO over a temperature range of 40–400 K at constant ambient pressure, following the same protocol as in Ref. 63. Once again, we recompute consistent energetics for the material-specific dataset, to build bespoke and finetuned PET models. Test-set MAEs on energy|forces are 12.67 meV/atom |27.96 meV/Å for PET-MAD, 0.23 meV/atom |9.41 meV/Å for the bespoke model, and 0.12 meV/atom |3.92 meV/Å for the LoRA-finetuned model, respectively.
@@
We then analyze the sampled configuration within a reduced-dimensionality collective variable landscape derived from a subset of neighbor density coefficients [63] (shown in the inset of Figure 9). Measuring the populations of different phases by a clustering algorithm allows us to estimate their relative chemical potential, and hence the transition temperatures between the different phases. Similar to what is observed in Ref. 63, the transition temperatures are heavily underestimated with respect to experiments (due to a combination of DFT shortcomings and finite-size effects), but there is excellent agreement between PET-MAD and the two bespoke models. The largest discrepancy, for the T-C transition, is below 30 K—a remarkable accuracy for a model with such broad applicability.
@@
To further assess the reliability of PET-MAD in predicting functional properties, we also compute the temperature-dependent static dielectric tensor $\boldsymbol{\epsilon}_{r}$ (Figure 9), by estimating the covariance of the total polarization, computed in turn using an equivariant linear model trained on the polarization dataset of Ref. 63. All models predict qualitatively the large value of $\boldsymbol{\epsilon}_{r}$ for the paraelectric, cubic phase, which increases greatly as it approaches the ferroelectric transition temperature, and then takes smaller values in the tetragonal and orthorhombic phases (which have multiple inequivalent optical axes). As it is the case of the transition temperatures, the two specialized models are in near-perfect agreement, while the PET-MAD simulations slightly underestimate $\boldsymbol{\epsilon}_{r}$ in the T phase, and yield a curve in the C phase that is shifted by about 25 K, consistent with the lower Curie temperature. Overall, this final example demonstrates once more that, despite considerably larger test errors than for a bespoke model, PET-MAD captures with close-to-quantitative accuracy subtle physical effects, and can be used to drive complicated, advanced materials modeling protocols.
@@ III Discussion
PET-MAD pushes the boundaries of what generally-applicable “universal” machine learning potentials can achieve, providing out-of-the-box semi-quantitative accuracy for both inorganic and organic materials benchmarks, as well as for six demonstrative applications covering a very broad range of material types and advanced simulation techniques. Moreover, it does so while being trained on a data set that is orders of magnitudes smaller than those used for current state-of-the-art models. This observation suggests that the computational budget for training set construction can be best allocated by pursuing internal consistency and a high degree of structural and chemical diversity — two guiding principles which underlie the construction of the MAD dataset we introduce here. A compact dataset facilitates model training and optimization, as well as the transition to more accurate, and computationally demanding, approximations to the electronic-structure problem.
The PET architecture we use is not constrained to follow rigorous rotational invariance; nonetheless, it learns to make predictions that are invariant to a high degree of accuracy. It achieves a high expressiveness thanks to its transformer module, while being faster than all the conservative models we considered, and demonstrating absolute speed records in the non-conservative regime. The availability of a reliable and inexpensive uncertainty quantification framework makes it possible to assess model error, and also its propagation to the ultimate quantity of interest. In cases where this error is larger than desired, a simple fine-tuning procedure can be applied to further improve the accuracy for a specific target system, while maintaining the ability to make qualitative predictions for the high-diversity test set. PET-MAD is easily accessible from several atomic-scale simulators thanks to its integration within the metatensor[64] ecosystem. It provides an efficient and accurate solution to perform simple and advanced atomistic simulations of materials, and we expect it will serve as inspiration for future developments of both datasets and model architectures, as well as a reliable engine to drive the design and discovery of new and better materials.
IV Methods
We provide a brief but complete overview of the procedure we followed to construct the MAD dataset, to train the PET-MAD model, and to fine-tune single-purpose models. The simulation protocols for the advanced simulation examples (Sections  II.4-II.9) follow closely those of the cited publications, and are briefly discussed in the SI for the convenience of the reader.
@@ IV.2 First-principles calculations
All datasets in this work (i.e. MAD and the reference subsets used for benchmarking, as well as the single-purpose dataset for the six case studies) were built by computing energies and forces using an identical density functional theory (DFT) setup. We used the Quantum Espresso v7.2 package [65], compiled with the SIRIUS[66] libraries as the primary DFT engine and AiiDA [67] as a workflow and task manager. To reduce the possibility that some simulations would converge to ambiguous magnetization states, all simulations were performed in a non-magnetic setting, using the PBEsol exchange-correlation functional. Even though this choice is known to have shortcomings for several classes of materials – as shown by the discrepancy of some of our examples with experiments – it allows us to generate predictions for a highly diverse system with a high degree of stability and converged numerical parameters. We described valence and semi-core electrons using the standard solid-state pseudopotentials library (SSSP) v1.2 (efficiency set) [68], with the most stringent settings for plane-wave and charge-density cutoffs across the 85 elements (110 Ry and 1320 Ry, respectively). As discussed in the SI, using different settings to match the maximum cutoff of each structure would give substantial inconsistencies between the description of some atomic types. Electronic smearing and partial occupancies were described with a Marzari-Vanderbilt-DeVita-Payne cold smearing with a spread of 0.01 Ry. In all periodic dimensions, the first Brillouin zone was sampled with a $\Gamma$-centered grid with a resolution of 0.125 Å-1, while only a $\Gamma$ point was used along the non-periodic dimensions. In the case of non-periodic systems we also applied a truncation scheme of the Coulomb potential to avoid the interaction of periodic replicas of the system through the periodic boundary conditions: the Sohier-Calandra-Mauri method for 2D systems and the Martyna-Tuckerman correction for 0D systems. Along each non-periodic direction we additionally applied a 25 Å vacuum to ensure convergence of the truncation methods. A compositional baseline based on isolated atom energies was subtracted from the DFT energies of each structure to improve the numerical stability of training.
For most of the MAD subsets listed in Section II.1, the DFT settings provided above ensured a good convergence rate of $95\%$, i.e. convergence was not achieved in less than 5% of the cases. The exception is MC3D-random, for which only about 55% of the attempted simulations converged. This result does not come as a surprise and can be explained by the extremely non-equilibrium state of these structures.
@@ IV.3 PET model
As our machine learning approach for fitting interatomic potentials, we employ the Point Edge Transformer (PET) [19], which has been shown to achieve state-of-the-art performance across a diverse set of benchmarks spanning both molecular and materials domains. In essence, PET is a graph neural network (GNN) where each message-passing layer is implemented as an arbitrarily deep transformer. Specifically, PET maintains feature vectors (or messages) $f_{ij}^{l}$ for every directed bond between atoms $i$ and $j$ that lie within a specified cutoff radius. Here, the superscript $l$ denotes the message-passing layer index, and each $f_{ij}^{l}$ is a fixed-size vector (a token) of dimension $d_{\mathrm{PET}}$. These intermediate representations are updated at each message-passing layer by a transformer with an adjustable number of internal layers, invoked for each atomic environment. At each message-passing layer, for each atom $i$, the input tokens to the transformer consist of the message vectors $\{f_{ji}^{l}\}_{j}$ coming from all neighbors $j$ to the central atom $i$. The transformer then applies a sequence-to-sequence transformation in a permutation-covariant manner. Its outputs are subsequently interpreted as the new set of outbound messages from atom $i$ to each neighbor $j$, $\{f_{ij}^{l+1}\}_{j}$. Geometric information (i.e., the 3D positions of neighbors) and chemical species are also incorporated into this process. To obtain the desired target property (e.g., energy), PET applies feed-forward neural networks to each representation $f_{ij}^{l}$ and sums all the outputs across bonds $ij$ and layers $l$. The PET architecture imposes no explicit rotational symmetry constraints, but learns to be equivariant through data augmentation. This unconstrained approach yields high theoretical expressivity: even a single layer of the model acts as a universal approximator featuring virtually unlimited body order and angular resolution. A detailed specification of the architecture and the full functional form can be found in Ref. 19.
@@ IV.4 Training of PET-MAD
We split the MAD dataset into training, validation, and test parts by separately shuffling each of the eight subsets, and selecting respectively 80%, 10%, and 10% of the structures. The composition contribution to the total energy of the structures was fitted using a simple linear model and subtracted to improve the training behavior. We used a Pareto-optimal architecture (see the details on hyperparameters optimization in the Supplementary Text) with a cutoff radius of 4.5 Å, 2 message-passing layers, each containing 2 transformer layers with a token size of 256, 8 heads in the multi-head attention layer, 512 neurons in the output multi-layer perceptron. Training was performed using the PyTorch framework and the metatrain package [64] on 8 NVIDIA H100 GPUs with a batch size of 24 structures per GPU for a total of 1500 epochs and took about 40 hours. We used the Adam optimizer with an initial learning rate (LR) of $10^{-4}$ and applied a LR scheduler, which halved the LR every 250 epochs. The loss function was based on a root-mean-square error difference between the model’s energy and force predictions and the corresponding target values, with a scaling factor of 0.1 applied to the energy contribution. Mean absolute errors of the trained PET-MAD model were 7.3 meV/atom|43.2 meV/Å  in energies and forces prediction on a train set. Validation set errors were 14.7 meV/atom|72.2 meV/Å  , respectively.
@@ IV.5 Fine-tuning
Fine-tuning of the pre-trained PET-MAD model was performed using a parameter-efficient fine-tuning technique based on low-rank adaption (LoRA)  [29]. This method is widely used in the ML community, since it allows to mitigate the catastrophic forgetting — the phenomenon of losing the accuracy on the base dataset while tuning the model on the new dataset, which is inherent to the conventional fine-tuning, where all the weights of the model are trainable [69]. In the LoRA approach, the weights of the base model are frozen during the training, while an additional set of trainable weights composed of two low-rank matrices is added to each attention block of the model with a regularization factor, which controls the influence of the low-rank matrices on the model’s weights. By adjusting rank and regularization factor, one can optimize the fine-tuning workflow to achieve better performance on specific tasks, without compromising entirely the ability to perform the more general task. In this work, we used a rank of 8 and a scaling parameter of 0.5 for all LoRA finetuned models, unless otherwise specified.
In terms of raw accuracy on the specialized datasets, the results in Section II and Section S10 indicate that fine-tuning is always beneficial compared to training a specialized ad-hoc model in the low-data regime. On larger datasets, while in some cases a specialized model trained from scratch can exceed the accuracy of a finetuned model, this is not the case for many of the systems investigated in this work, including barium titanate, succinic acid, lithium thiophosphate, and high-entropy alloys.
LoRA-finetuned models retain varying degree of accuracy (see the Section S9 for details) on the generic structures from the MAD dataset, while still providing practically equivalent observables from an ad-hoc fully specialized model. In general, we therefore recommend to employ LoRA if finetuning PET-MAD for a specific application.